SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (6th Feb, 2007)

Attendees: Debbie Bard, Brandon Drummond, Al Eisner, Hojeong Kim, Selina Li, Rahmat Rahmat, Su Dong, Mark Tibbetts.

Operations

  • During the ROD last wednesday the spae BLT and GLT boards were swapped in and cosmic runs taken. The spare GLT board caused high deadtime and damage (see log files here at 17.12) and the spare BLT board caused some funny looking distributions. It's possible the FCT PROM that was replaced in the spare GLT was incorrect, and any errors in that wouldn't show up at the teststand unless you did a DAQ test at the teststand. The BLT board may not have been the correct spare, or there may have been correction errors. The nect scheduled ROD is in ~3 weeks, and we ought to have the boards ready to test again by then.
  • The super new official on-call trigger cellphone number is 650-238-5824.

EMT

  • Still not spoken to Jong about the problems he mentionned.
  • The EMC has a couple more dead channels, not causing any problems in the EMT.
  • Threshold study: The macros to read the L1TOprMonXtcApp hbook files are in fortran. They're being converted to root, but should run the macros over old runs anyway to have an idea of what to expect.
  • The error in the EmtOprMon module has been fixed (by person or persons unknown). In looking at the trackmatching, Mark noticed it's not used at all - see the end of this. Something nice could be added here.

L3

  • It appears the tracking group forgot about us! They have been reminded that the changes in the L3 tracking need validating.

Data Quality and Monitoring

  • The pi0 yield in Run 6 so far is about 20% higher than for Run 5. It's not yet clear whether this is an effect of the high background conditions (more combinatoric pi0s) or a real increase in the number of pi0s (more pi0s per hadronic event).
  • See sample stripchart for example with lines between points (Thanks to Hojeong for the help). There is also auto-scaling of the y-axis - it would be best to keep the lines and have a fixed scale instead, so the errors can be seen clearly and it's obvious when a point goees off the scale.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.